OpenStudy (anonymous):

MacLaurin series cos(x^2)

OpenStudy (anonymous):

Supposed to find the first 4 non-zero terms of the series above

OpenStudy (anonymous):

I get: f(0)=1 f'(0)=0 f''(0)=0 f'''(0)=0

OpenStudy (anonymous):

wondering if its ever gonna not be zero besides the first term...

OpenStudy (anonymous):

f(x) = cos(x)^2 f'(x) = -2 cos(x) sin(x) f''(x) = 2 sin^2(x) - 2 cos^2(x) f'''(x) = 8 cos(x) sin(x) f''''(x) = -8 sin(x) + 8 cos(x) f'''''(x) = -32 cos(x) sin(x) f''''''(x) = 32 sin^2(x) - 32 cos^2(x) f(0) = 1 f'(0) = 0 f''(0) = -2 f'''(0) = 0 f''''(0) = 8 f'''''(0) = 0 f''''''(0) = -32 Is this of any help?

OpenStudy (anonymous):

I think remnant is true

OpenStudy (anonymous):

Not (cos(x))^2 but cos(x^2)

OpenStudy (anonymous):

please read question properly: cos(x^2)

OpenStudy (anonymous):

So for the first four series the answer is 1

OpenStudy (anonymous):

i got:\[f^{(4)}(0)=-8\] so i guess it does become non zero

OpenStudy (anonymous):

im sorry but it looks like everyone else has got this so i will go elsewhere

OpenStudy (anonymous):

\[\Large{ \cos(x^2) = \sum_{k=0}^\infty \frac{(-1)^k (x^2)^{2 k}}{(2 k)!}\\=1-\frac{x^4}{2}+\frac{x^8}{24}-\frac{x^{12}}{720}+\cdots}\]

OpenStudy (anonymous):

just replace x from the maclaurin series for cos x as x^2, wallah you are done, no differentiation needed